{2x+y=5
{x^2+y2=50
Solve the nonlinear system by substitution

Answer by Theo(13342)

solve for y in the first equation to get y = 5 - 2x

replace y in the second equation with 5 - 2x to get:

x^2 + (5-2x)^2 = 50

simplify to get:

x^2 + 25 - 20x + 4x^2 = 50

combine like terms and reorder the terms to get:

5x^2 - 20x + 25 = 50

subtract 50 from both sides of the equation to get:

5x^2 - 20x - 25 = 0

factor to get:

(5x+5) * (x-5) = 0

solve for x to get:

x = -1 and x = 5

in either equation, solve for y to get:

when x = -1, y = 7

when x = 5, y = -5

when you solve for y in the first equation, you will get an ambiguous answer.

for example:

x^2 + y^2 = 50

when x = -1, you get y = plus or minus sqrt(49).

when x = 5, you get y = plus or minus sqrt(25)

y could be -7 or plus 7 when x = -1 based on the first equation.
y could be -5 or plus 5 when x = 5 based on the first equation.

however, the second equation tells you that:

when x = 5, y = 5 - 2x = 5 - 10 = -5
when x = -1, y = 5 - 2(-1) = 5 + 2 = 7

there is no ambiguity there and since the solution needs to satisfy both equations, your answer is:

(5,-5) is one of the coordinate points common to both equations.
(-1,7) is the other of the coordinate points common to both equations.

the graph of both equations is shown below:

$$$